OpenStudy (anonymous):

200=1/32(100)^2 sin(2x) how do i get the value for x?

OpenStudy (anonymous):

\[200=\frac{1}{32}(100)^2\sin(2x)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x is theta

OpenStudy (anonymous):

\[6400=(100)^2\sin(2x)\] \[.64=\sin(2x)\] \[2x=\sin^{-1}(.64)\] and a calculator